Find the area of the triangle formed from the coordinate axes and the tangent line to the curve y = 5x^(−1) −x/5 at the point (5,0).


1
Expert's answer
2021-10-18T07:47:28-0400

Let us find the area of the triangle formed from the coordinate axes and the tangent line to the curve "y = \frac{5}{x}−\frac{x}5" at the point (5,0). It follows that "y' = -\frac{5}{x^2}−\frac{1}5," and hence "y'(5)=-\frac{2}5." The equation of the tangent to the curve at the point (5, 0) is "y=-\frac{2}5(x-5)." If "x=0," then "y=2," and hence the point (0, 2) is the y-intercept. We conclude that the area of the triangle is equal to "\frac{1}2\cdot 5\cdot 2=5" squared units.